OpenStudy (anonymous):

A line passes through the origin and has a slope of –3. What is the equation of the line that is perpendicular to the first line and passes through the point (3, 4)?

OpenStudy (texaschic101):

a line that passes through the origin (0,0) and has a slope(m) of -3 y = mx + b 0 = -3(0) + b 0 = 0 + b 0 = b so your equation is : y = -3x + 0 however, we are looking for perpendicular. That means that the slopes have to be negative reciprocals of each other. All that means is " flip " the slope and change the sign. So the negative reciprocal of -3 is 1/3 (see how I flipped the slope and changed the sign) y = mx + b slope(m) = 1/3 (3,4)...x = 3 and y = 4 now we sub 4 = 1/3(3) + b 4 = 1 + b 4 - 1 = b 3 = b so your perpendicular equation is : y = 1/3x + 3

OpenStudy (anonymous):

thanks!